OpenStudy (anonymous):

can someone please help & explain this question to me ? Which of the following are measurements of the sides of a right triangle? A.10,8,6 B.13,12,5 C.26,24,10 D.all of the above.

OpenStudy (asnaseer):

Do you know which property is satisfied by all right angled triangles?

OpenStudy (anonymous):

like do you mean what all the angles add up to be ?

OpenStudy (asnaseer):

no - have you heard of the Pythagorus Theorem?

OpenStudy (anonymous):

yes i have a^2+b^2=c^2 .

OpenStudy (asnaseer):

great! that is what you need to use here. in that equation, 'c' represents the hypotenuse of the triangle - which is the longest side. so what you need to do is to check which of the triples A, B, or C satisfy this equation.

OpenStudy (asnaseer):

so start with A, does this equation hold true?\[8^2+6^2=10^2\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good, now try the same check with the triples in B and C

OpenStudy (anonymous):

oH ! okay so now i get it ! it wasnt so hard after all ! thanks a million ! (:

OpenStudy (asnaseer):

yw :)

OpenStudy (asnaseer):

BTW: what answer did you get?

OpenStudy (anonymous):

i got answer 'A' ! b and c were really weird answers .

OpenStudy (asnaseer):

please double check your work - remember 'c' represents the longest side, so the equations you need to check are:\[12^2+5^2=13^2\]and:\[10^2+24^2=26^2\]

OpenStudy (anonymous):

wait im confused right now .

OpenStudy (asnaseer):

in what part are you confused?

OpenStudy (anonymous):

like im working out the equation and like no of them seem to work

OpenStudy (asnaseer):

ok, lets take the triples in B first, which were 12, 5 and 13 now what is \(12^2\)?

OpenStudy (anonymous):

144

OpenStudy (asnaseer):

correct, and what is \(5^2\)?

OpenStudy (anonymous):

25

OpenStudy (asnaseer):

correct again, so we get:\[12^2+5^2=144+25=169\]does this equal \(13^2\)?

OpenStudy (anonymous):

yes it does

OpenStudy (asnaseer):

good, so now we know that both A and B satisfy the Pythagorus Theorem. now lets try the triples in C which were: 10, 24 and 26. what is \(10^2+24^2\)?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

soo its all of them ?

OpenStudy (asnaseer):

yes :)

OpenStudy (asnaseer):

a time saving trick you can use with these triples is to first divide all the triples with any common factors that they may have

OpenStudy (asnaseer):

e.g. In A) 10, 8 and 6 can all be divided by 2, so we end up with: 5, 4 and 3 which means we just need to check if:\[3^2+4^2=5^2\]

OpenStudy (anonymous):

ooo okay i will be sure to use that from now on ! now i really get it, i was using the wrong numbers the first time !

OpenStudy (asnaseer):

ok good - well done! :)

OpenStudy (anonymous):

again THANK YOU !(:

OpenStudy (asnaseer):

again - yw :)